# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v2

@@ Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data. In addition, we set new benchmarks for two foundational speech decoding tasks. Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data for speech decoding. We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled neuroimaging data, permitting learning from large heterogeneous datasets. The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation. We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
@@ 1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases [50]. In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains [29, 7, 42, 46].
@@
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present (e.g. [17]), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g. [27]), holds the promise of training models on collective, internet-scale data.
@@
Given the scarcity of labelled data, self-supervised learning (SSL) appears promising as an avenue for domains where such data is rare or hard to obtain [3]. In the SSL paradigm, pretext tasks pre-train a model on unlabelled data by generating implicit training labels through transformations of the input data in order to help a downstream task. We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals. In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive magnetoencephalography (MEG), setting the baselines for speech detection and voicing classification on this data. The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time. Our main contributions are:
@@
  • A set of domain specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
@@
  • A neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
  • A comprehensive experimental evaluation providing evidence that our approach can scale up with data and enable cross-dataset, task, and subject generalisation.
@@ 2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies [40, 39, 60, 9] and non-invasive studies [52]. These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Focusing on semantic rather than phonetic decoding, work by Tang et al. [52] is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos. They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions. Csaky et al. [11] find that a subset of individuals perform better when evaluated with a group-level model than with individual models. Exploiting audio data in a multi-modal framework, Défossez et al. [17] show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated. Although they repeat the result within two MEG and two EEG datasets, Défossez et al. [17] do not show any improvements for pooling data across datasets. Moreover, they do not combine data from studies with different labels either; cf. [58, 16, 56]. Unfortunately, two of these papers [58, 16] included a bug in their evaluation code. As such, their methods may perform no better than a baseline that provides pure noise inputs to the model [28].
@@
In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning [1, 15, 41, 31, 62, 21]. Until now, similar approaches have not translated to the brain decoding literature. However, prior work has used other methods to leverage unlabelled brain data. For example, Jiang et al. [27] succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective. Although this work combined unlabelled EEG datasets, it only achieved improvements on non-speech tasks. Wang et al. [57] used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations. They attained impressive speech detection results but with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
@@ 3 Method
To process continuous neuroimaging data, we introduce a neural architecture for encoding heterogeneous brain signals into latent representations. By developing pretext tasks with the objective of learning generalisable brain representations, we leverage this architecture for self-supervised learning from unlabelled data, hoping to replicate similar successes in computer vision [21, 8].
@@ 3.1 Network Architecture
We design a two-stage neural network architecture (Figure 2). The pre-training stage uses pretext tasks to train a representation with unlabelled brain data. Then, the fine-tuning stage uses this representation to learn the downstream task by fine-tuning with labelled data.
@@
Figure 2: Architecture overview. Inputs are projected into a shared dimension, then encoded. In pre-training, all weights are trainable except for modules in light-red, while in shallow fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen. Deep fine-tuning is identical, except the encoder is trainable. Dashed borders indicate optional components.
Normalisation. We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance.
@@
Backbone. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs [61, 14]. Specifically, our convolutional encoder adapts the implementation of the SEANet architecture [51] used in Défossez et al. [14]. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
Pre-training. Following the advice of Balestriero et al. [3, Section 3.2], we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks.
Fine-tuning. In this stage, we introduce classifiers for the downstream tasks and train with labelled data and supervised learning (Figure 2). Depending on the experiment, we fine-tune in one of two ways. In the first, we train an MLP classifier from scratch on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. We call this shallow fine-tuning. In the other, we fine-tune by training an MLP on top of the pre-trained representation again, but do not freeze the backbone, training end-to-end. Here, we backpropagate through the classifier and the backbone. We refer to this as deep fine-tuning. In both cases, a new trainable dataset-specific linear layer can be introduced for a novel dataset.
For speech detection, our classifier makes a prediction for each individual embedding. For voicing classification, where there is only one label for each sample window, the embeddings are flattened into a tensor in $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ representing the entire window. This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature [12].
@@
Subject conditioning. Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects [11]. In the speech literature, models include speaker conditioning to account for these differences [20]. We take a similar approach by introducing subject conditioning, represented as a subject embedding, to our model. With a SEANet-based architecture, Zeghidour et al. [61] find that conditioning is equally as effective at the encoder bottleneck as in other stages of the model. Hence, we place ours at the cortex encoder bottleneck for simplicity. We evaluate multiple types of conditioning including subject embeddings and feature-wise linear modulation (FiLM) [44].
@@ 3.2 Pretext Tasks
Our pretext tasks are unsupervised feature learning tasks for continuous neuroimaging signals that aim to learn generalisable speech decoding features. Since different datasets use different hardware and varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal. This means that these tasks do not require identifying specific sensor channels.
Band prediction. In the literature, neural responses can be broadly segmented into functional frequency bands [22, 45, 36]. Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech [35], Theta ($\theta$) waves (4–8 Hz) reliably track [34] and phase-lock to the amplitude envelope of heard sentences [43], Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals [49], Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding [5] which affects lexical processing [59], Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g. memory, learning, reasoning, and planning) [19, 6], and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection [26] and phonemic feature classification in the STG [38] as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) [10]. As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected. This yields the loss
@@
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity [18, 54]. Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work [27] also finds phase to be an essential component of the signal.
@@
Amplitude scale prediction. MEG and EEG signals use an array of sensors at different spatial locations, capturing different signal sources more intensely. Representing the relative amplitude difference between sensors could be important for differentiating between neural responses originating from distinct parts of the brain. Within speech, Hamilton et al. [26] find that localised regions of the STG respond to sustained speech and speech onsets. Differentiating between neural responses from this region and others may be essential for decoding speech perception.

### v3

@@ Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit individual differences, such as anatomy, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. This gap exists for all neural data, but especially for magnetoencephalography (MEG), where the scale of individual datasets has not yet caught up with other modalities. To address this, we develop a set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results with MEG show that representations learned with these objectives scale with data, generalise across subjects, datasets, and tasks, outperform using the raw input representation, and even surpass comparable self-supervised approaches. In addition, we set new benchmarks for two foundational speech decoding tasks. Collectively, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data using pretext tasks for speech decoding. We pre-train a neural network using tasks that generate implicit labels from abundant unlabelled MEG neuroimaging data, permitting learning from large heterogeneous datasets. The tasks apply a randomly selected neuroscientifically relevant transformation $T$ to the data and the network predicts the transformation. We then train a linear probe on top of the pre-trained model, which remains frozen, with labelled data, achieving superior generalisation (cf. raw inputs) owing to the strength of the representation.
@@ 1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases (Sutton, 2019). In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains (Jumper et al., 2021; Caron et al., 2021; OpenAI, 2023; Radford et al., 2023).
@@
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present, e.g. Défossez et al. (2023), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields, such as  Jiang et al. (2024), holds the promise of training models on collective, internet-scale data.
@@
While neuroimaging modalities such as electroencephalography (EEG) are more abundant, MEG may be a better modality for decoding as it provides a richer signal (Lopes da Silva, 2013; Hall et al., 2014). Given the scarcity of speech-labelled MEG data and the relative abundance of other MEG data, self-supervised learning (SSL) appears promising as it is an avenue for domains where labels are rare or hard to obtain (Balestriero et al., 2023). But the scale of public MEG data, while large, is still not at the volume of breakthroughs in self-supervised image and natural language processing, let alone EEG. Thus, SSL methods for MEG need to be highly data-efficient. Pretext tasks are one such method in which domain-specific self-supervised tasks are used to pre-train a model on unlabelled data by generating implicit training labels through transformations of the input in order to help a downstream task. We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals which we train using our pretext tasks. In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive MEG, setting the baselines for speech detection and voicing classification on this data. The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time. Our main contributions are:
@@
  • A set of domain-specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
@@
  • A data-efficient neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
  • A comprehensive experimental evaluation, using multiple times the volume of data in prior work, that verifies the above claims and additionally provides evidence for the existence of scaling laws when pre-training models with unlabelled MEG recordings.
@@ 2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies (Moses et al., 2021; Metzger et al., 2023; Willett et al., 2023; Chen et al., 2024) and non-invasive studies (Tang et al., 2023). These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Focusing on semantic rather than phonetic decoding, the method developed by Tang et al. (2023) is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos. They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions. Csaky et al. (2022) find that a subset of individuals perform better when evaluated with a group-level model than with individual models. Exploiting audio data in a multi-modal framework, Défossez et al. (2023) show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated. However, they do not demonstrate the ability to generalise to novel subjects and must retrain their model for new datasets. Moreover, although they repeat the result within two MEG and two EEG datasets, Défossez et al. (2023) do not show any improvements for pooling data across datasets. Their method is also unable to incorporate data without corresponding audio labels and so they do not combine data from studies with other kinds of labels either; cf. Wang & Ji (2022); Duan et al. (2023); Wang et al. (2023a). Unfortunately, the first two of these papers included a bug in their evaluation code. As such, their methods may perform no better than a baseline that provides pure noise inputs to the model (Jo et al., 2024).
@@
In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning (Agrawal et al., 2015; Doersch et al., 2015; Noroozi & Favaro, 2016; Larsson et al., 2016; Zhang et al., 2016; Gidaris et al., 2018). Until now, similar approaches have not translated to the brain decoding literature. However, prior work has used other methods to leverage unlabelled brain data. For example, Jiang et al. (2024) succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective. Although this work combined unlabelled datasets, their results studied simpler non-speech tasks with EEG. Wang et al. (2023b) used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations. Their impressive speech detection results were achieved with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data. Perhaps the closest work to ours in terms of unlocking scaling with neural data is BIOT (Yang et al., 2023). This is a self-supervised architecture for encoding bio-signals that is similarly capable of training with different datasets, labels, and varied numbers of sensors. Like the previous works, the approach tokenises signals for a transformer architecture, but instead of a masked loss it uses a contrastive pre-training objective. While theoretically supporting MEG, Yang et al. (2023) evaluate BIOT on simple ECG/EEG tasks rather than address the comparatively complex challenge of speech decoding with MEG data.
@@ 3 Method
To encode continuous neuroimaging data, we introduce a neural architecture to embed heterogeneous brain signals. We leverage this architecture for self-supervised learning from unlabelled MEG data using a set of pretext tasks designed to generate generalisable brain representations for speech decoding. With this approach, we hope to replicate similar successes in computer vision (Gidaris et al., 2018; Chen et al., 2020).
@@ 3.1 Network Architecture
Our two-stage neural network architecture (Figure 2) uses pretext tasks in pre-training to learn a representation with unlabelled brain data. Then, the fine-tuning stage uses this representation to learn the downstream task by training with labelled data.
@@
Figure 2: Architecture overview. Inputs are projected into a shared dimension by the dataset-conditional layer, then encoded. In pre-training, all weights are trainable except for modules in light-red, while in fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen. Dashed borders indicate optional components.
@@
We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs (Zeghidour et al., 2022; Défossez et al., 2022). Specifically, our convolutional encoder adapts the implementation of the SEANet architecture (Tagliasacchi et al., 2020) used in Défossez et al. (2022) which we describe here and as part of Figure 2. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
@@
Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects (Csaky et al., 2022). In the speech literature, models include speaker conditioning to account for these differences (Gibiansky et al., 2017). We take a similar approach by introducing subject conditioning. Zeghidour et al. (2022) find that conditioning is equally effective at the encoder bottleneck as in other stages of the model. Hence, we place ours at the cortex encoder bottleneck for simplicity. We use feature-wise linear modulation (FiLM) (Perez et al., 2018) as our conditioning method.
Following the advice of Balestriero et al. (2023, Section 3.2), we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks. When fine-tuning, we train a linear decoder, for a downstream task, on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. A trainable dataset-specific linear layer can be introduced for a novel dataset.
For speech detection, our classifier makes a prediction for each individual embedding. For voicing classification, where there is only one label for each sample window, the embeddings are flattened into a tensor in $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ representing the entire window. This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature (Csaky et al., 2023).
@@ 3.2 Pretext Tasks
Our pretext tasks are unsupervised feature learning tasks that aim to learn generalisable speech decoding features. Since different datasets use varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal.
Band prediction. In the literature, neural responses can be segmented into functional frequency bands (Giraud & Poeppel, 2012; Piai et al., 2014; Mai et al., 2016). Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech (Luo et al., 2010), Theta ($\theta$) waves (4–8 Hz) reliably track (Luo & Poeppel, 2007) and phase-lock to the amplitude envelope of heard sentences (Peelle et al., 2012), Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals (Strauß et al., 2015), Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding (Bressler & Richter, 2015) which affects lexical processing (Weiss & Mueller, 2012), Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g. memory, learning, reasoning, and planning) (Fries, 2009; Buzsáki & Wang, 2012), and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection (Hamilton et al., 2018) and phonemic feature classification in the STG (Mesgarani et al., 2014) as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) (Cheung et al., 2016). As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding linear predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected. This yields the loss
@@
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity (Fries, 2005; Vidaurre et al., 2018). Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work (Jiang et al., 2024) also finds phase to be an essential component of the signal.
@@
Amplitude scale prediction. MEG and EEG signals use an array of sensors at different spatial locations, capturing different signal sources more intensely. Representing the relative amplitude difference between sensors could be important for differentiating between neural responses originating from distinct parts of the brain. Within speech, Hamilton et al. (2018) find that localised regions of the STG respond to sustained speech and speech onsets. Differentiating between neural responses from this region and others may be essential for decoding speech perception.